# General Oscillator-Based Ising Machine Models with Phase-Amplitude Dynamics and Polynomial Interactions

### v2

@@ I INTRODUCTION
The limitations of traditional von Neumann architectures in efficiently tackling combinatorial optimization problems have spurred the exploration of alternative computing paradigms. Among these, Ising machines, inspired by the Ising model from statistical mechanics, have emerged as a promising approach. The versatility of Ising machines extends across a diverse range of applications, including portfolio optimization [1, 2], drug discovery [3], vehicle routing [4, 5, 6], fraud detection [7], clustering [8], power grid optimization [9], wireless communication [10, 11], image restoration [12] and machine learning [13, 14, 15, 16, 17]. Additionally, Ising machines leverage multiple specialized hardware substrates, including quantum [18, 19], optical [20, 21, 22, 23, 24, 25, 26, 27, 28, 29], CMOS [30, 31, 32, 33, 34, 35, 36, 37] and other emerging hardware platforms [38, 39, 40, 41, 42, 43, 44]. By mapping the decision variables of an optimization problem onto interacting spins [45], Ising machines can efficiently search for solutions by allowing the system to naturally evolve towards its lowest energy state. This process corresponds to finding a configuration of spins that minimizes the system’s energy, which directly translates to a high-quality solution to the original optimization problem. This approach offers the potential to significantly accelerate the solution of complex optimization problems that are intractable for classical computers.
@@
Ising machines based on networks implementing the Kuramoto model [55] of synchronization have gained significant traction. These oscillator-based Ising machines (OIMs) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] based on Kuramoto model exploit the emergent collective behavior of synchronized oscillators to emulate spin interactions, with the system’s natural tendency towards stable, synchronized states corresponding to low-energy solutions of the Ising Hamiltonian. While Kuramoto-based OIMs have shown promise, oscillator models that incorporate amplitude dynamics in addition to phase [51, 68, 69, 70, 71] have aimed to provide a more complete physical description of the underlying hardware. Research has revealed that phase-amplitude formulations for nanoelectromechanical (NEMS) oscillators can diverge significantly from simplified phase-only descriptions under specific operating regimes [68]. Moreover, incorporating amplitude dynamics into spintronic oscillator models (Slavin’s model [69]) yields a more faithful representation of coupled magnetization processes and the system demonstrates improved time complexity over the Goemans-Williamson algorithm [71]. For consistency with [51], we refer to models that integrate both amplitude and phase in the context of Ising machines as the Hopf oscillator model.
@@
To address these limitations, we have developed an oscillator model with complex-valued oscillator representation, based on Andronov-Hopf dynamics [72], providing a rigorous framework for oscillator-based Ising machines. Our key contributions include:
@@
This rigorous framework not only provides a deeper theoretical understanding of OIMs but also leads to significant empirical performance improvements compared to existing phase-amplitude approaches. Remarkably, the energy function terms in our model closely resemble those found in nonlinear optical or four-wave mixing Hamiltonian [74, 75]. This connection suggests exciting new paths for realizing powerful Ising machines, potentially leveraging existing photonic technologies. Furthermore, we demonstrate that our model reduces to a previously proposed high-order, Kuramoto-based formulation [54] once amplitude dynamics are removed. Our work therefore generalizes previous polynomial frameworks, offering perspective on the relationship between Hopf and Kuramoto oscillator models for Ising machines.
@@ II.1 Ising Hamiltonian for Combinatorial Optimization
1-D spin glass systems were originally proposed as a model for quenched disorder and frustration in metallic alloys [76, 77]. For $N$ 1-D spins $s=(s_{1},...,s_{N})$ with a coupling matrix $J\in\mathbb{R}^{N\times N}$ and external bias $h\in\mathbb{R}^{N}$, the Ising spin glass (ISG) Hamiltonian, named after Ernst Ising for his earlier work on purely ferromagnetic systems [78], is given by
@@
Later work revealed that finding the ground state of a spin glass is NP-hard [79, 80]. Accordingly, a number of high-impact combinatorial problems can be mapped to an equivalent spin glass system [45, 81]. In particular, an ISG Hamiltonian can be trivially mapped to Quadratic Unconstrained Binary Optimization (QUBO) problems, a well-studied class within operations research [82].
The relative simplicity of the ISG Hamiltonian makes it highly amenable for hardware acceleration, leading to the advent of “Ising machines” (IMs) [81]: devices which “implement” the Ising Hamiltonian in some manner. IM proposals include purely digital accelerators [83, 84], analog devices evolving a continuous relaxation of the Ising Hamiltonian [56, 35, 85], and hybrid digital/analog systems [86, 87].
@@ II.2 Solving 3-SAT problems with Ising Machine
Ising machines have been successfully applied to quadratic unconstrained binary optimization (QUBO) problems [35, 56, 88] and even polynomial unconstrained binary optimization (PUBO) [54, 89, 51] problems with the help of higher-order interactions. A generic Ising machine workflow usually involves ① converting the problem to a discrete Ising formula, ② mapping the formula to a continuous-time, dynamical Ising machine capable of seeking the minimum energy state and ③ extracting the solution from the final device state. In this subsection, we demonstrate how to map 3-SAT problems to an Ising Machine as a preliminary step.
@@
where each $\ell_{j,1}$ is a literal (either a variable $x_{p}$ or its negation $\overline{x}_{p}$) and each disjunction of literals $(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is a clause. The 3-SAT problem consists of finding some $x$ such that $f(x)=1$, i.e., an assignment satisfying all of the clauses. We restrict ourselves to 3-SAT without loss of generality, as any $K$-SAT problem can be reduced to an equivalent 3-SAT problem by adding $O(K\cdot m)$ variables and clauses [90].
@@ II.3 Kuramoto Oscillator Model
where $\kappa$ and $J_{ij}$ describe global and pairwise interaction strengths respectively, and $\omega_{i}$ is the natural frequency of the $i^{th}$ oscillator. We set $\omega_{i}=0$ in Eq. 8 assuming that all oscillators have identical $\omega_{i}$ and viewing the system in a rotating reference frame. Using spin glass interaction coefficients for $J$ has led to oscillator-based accelerators for combinatorial optimization: oscillator Ising machines (OIM) [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 93, 94].
@@
The coupled oscillator system tends to minimize energy $E(\theta)$ over time [95], laying the foundation for OIMs that tackle combinatorial optimization problems. The original OIM architecture [56] utilized electronic LC-tank oscillators, and more recent proposals have broadened the choice of substrate to ring oscillators [96, 97], phase transition nano-oscillators [98, 99, 100], and magnetic tunnel junctions [101, 102].
@@ III.1 Hopf Oscillator Model
While the Kuramoto model is sufficient to study phase-only phenomena, real-world oscillators will have variability in both phase and amplitude. To enable a more faithful model, we can generalize from the phase-reduced representation to complex $z\in\mathbb{C}^{N}$. An oscillator model for combinatorial optimization and sampling is the Adronov-Hopf model (or “Hopf model” for brevity) [51, 72]. The evolution of each Hopf oscillator is given by a system of coupled ODEs
@@ IV.1 Model Evaluation
Overall Performance: We evaluate model performance on 1,700 uniform random 3-SAT benchmarks from SATLIB [103]. Each set, ufN-M, denotes instances with N variables and M clauses; we use uf20-91, uf50-218, uf75-325, uf100-430, and uf150-645, all at a clause-to-variable ratio of approximately 4.3. This ratio lies at the empirical phase transition in satisfiability, where random 3-SAT instances shift from mostly satisfiable to mostly unsatisfiable and where difficulty peaks [104]. We define a solvable instance as one where at least one of 100 independent runs (with different initial conditions) reaches zero energy within a sufficient simulation time ($T=136$), allowing both models to achieve their best performance on the benchmark.
@@ ACKNOWLEDGMENTS
The data that support the findings of this article are openly available [106].

### v3

@@ I INTRODUCTION
The limitations of traditional von Neumann architectures in efficiently tackling combinatorial optimization problems have spurred the exploration of alternative computing paradigms. Among these, Ising machines, inspired by the Ising model from statistical mechanics, have emerged as a promising approach. The versatility of Ising machines extends across a diverse range of applications, including portfolio optimization [tanahashi2019application, parizy2022cardinality], drug discovery [mao2023chemical], vehicle routing [bao2023ising, tsuyumine2024optimization, azad2022solving], fraud detection [stein2023exploring], clustering [matsumoto2022distance], power grid optimization [kirihara2023exploring], wireless communication [singh2022ising, sreedhara2023mu], image restoration [6621708] and machine learning [niazi2024training, wuextending, song2024ds, instatrain, liuexpressive]. Additionally, Ising machines leverage multiple specialized hardware substrates, including quantum [johnson2011quantum, king2023quantum], optical [clements2017gaussian, pierangeli2019large, marandi2014network, wu2025monolithically, calvanese2021all, wang2013coherent, cen2022large, gao2024photonic, pierangeli2020noise, sun2022quadrature], CMOS [takemoto20192, matsubara2020digital, belletti2008janus, baity2014janus, yamaoka201520k, 1, su2024flexspin, xie2022ising] and other emerging hardware platforms [honjo2021100, litvinenko2023spinwave, yue2024scalable, chen2024oscillatory, maher2024cmos, si2024energy, chen2025off]. By mapping the decision variables of an optimization problem onto interacting spins [21], Ising machines can efficiently search for solutions by allowing the system to naturally evolve towards its lowest energy state. This process corresponds to finding a configuration of spins that minimizes the system’s energy, which directly translates to a high-quality solution to the original optimization problem. This approach offers the potential to significantly accelerate the solution of complex optimization problems that are intractable for classical computers.
@@
Ising machines based on networks implementing the Kuramoto model [19] of synchronization have gained significant traction. These oscillator-based Ising machines (OIMs) [32, 4, 11, 24, 2, 7, graber2024firefly, sreedhara2024novel, garg2024phase, ochs2021ising, sreedhara2023digital, cilasun2025coupled] based on Kuramoto model exploit the emergent collective behavior of synchronized oscillators to emulate spin interactions, with the system’s natural tendency towards stable, synchronized states corresponding to low-energy solutions of the Ising Hamiltonian. While Kuramoto-based OIMs have shown promise, oscillator models that incorporate amplitude dynamics in addition to phase [6, matheny2019exotic, slavin2009nonlinear, gambuzza2016amplitude, garg2024improved] have aimed to provide a more complete physical description of the underlying hardware. Research has revealed that phase-amplitude formulations for nanoelectromechanical (NEMS) oscillators can diverge significantly from simplified phase-only descriptions under specific operating regimes [matheny2019exotic]. Moreover, incorporating amplitude dynamics into spintronic oscillator models (Slavin’s model [slavin2009nonlinear]) yields a more faithful representation of coupled magnetization processes and the system demonstrates improved time complexity over the Goemans-Williamson algorithm [garg2024improved]. For consistency with [6], we refer to models that integrate both amplitude and phase in the context of Ising machines as the Hopf oscillator model.
@@
To address these limitations, we have developed an oscillator model with complex-valued oscillator representation, based on Andronov-Hopf dynamics [nikonov2015coupled], providing a rigorous framework for oscillator-based Ising machines. Our key contributions include:
@@
This rigorous framework not only provides a deeper theoretical understanding of OIMs but also leads to significant empirical performance improvements compared to existing phase-amplitude approaches. Remarkably, the energy function terms in our model closely resemble those found in nonlinear optical or four-wave mixing Hamiltonian [Kanao_2021, quinn2024coherentisingmachinebased]. This connection suggests exciting new paths for realizing powerful Ising machines, potentially leveraging existing photonic technologies. Furthermore, we demonstrate that our model reduces to a previously proposed high-order, Kuramoto-based formulation [5] once amplitude dynamics are removed. Our work therefore generalizes previous polynomial frameworks, offering perspective on the relationship between Hopf and Kuramoto oscillator models for Ising machines.
@@ II.1 Ising Hamiltonian for Combinatorial Optimization
1-D spin glass systems were originally proposed as a model for quenched disorder and frustration in metallic alloys [edwards_theory_1975, sherrington_solvable_1975]. For $N$ 1-D spins $s=(s_{1},...,s_{N})$ with a coupling matrix $J\in\mathbb{R}^{N\times N}$ and external bias $h\in\mathbb{R}^{N}$, the Ising spin glass (ISG) Hamiltonian, named after Ernst Ising for his earlier work on purely ferromagnetic systems [ising1925beitrag], is given by
@@
Later work revealed that finding the ground state of a spin glass is NP-hard [barahona_computational_1982-2, istrailStatisticalMechanicsThreedimensionality2000]. Accordingly, a number of high-impact combinatorial problems can be mapped to an equivalent spin glass system [21, mohseni_ising_2022-1]. In particular, an ISG Hamiltonian can be trivially mapped to Quadratic Unconstrained Binary Optimization (QUBO) problems, a well-studied class within operations research [beasley_or-library_1990].
The relative simplicity of the ISG Hamiltonian makes it highly amenable for hardware acceleration, leading to the advent of “Ising machines” (IMs) [mohseni_ising_2022-1]: devices which “implement” the Ising Hamiltonian in some manner. IM proposals include purely digital accelerators [aramon_physics-inspired_2019, goto_high-performance_2021], analog devices evolving a continuous relaxation of the Ising Hamiltonian [32, 1, rajak_quantum_2023], and hybrid digital/analog systems [inagaki_coherent_2016, hizzani_memristor-based_2024].
@@ II.2 Solving 3-SAT problems with Ising Machine
Ising machines have been successfully applied to quadratic unconstrained binary optimization (QUBO) problems [1, 32, honjo_100000-spin_2021-1] and even polynomial unconstrained binary optimization (PUBO) [5, 29, 6] problems with the help of higher-order interactions. A generic Ising machine workflow usually involves ① converting the problem to a discrete Ising formula, ② mapping the formula to a continuous-time, dynamical Ising machine capable of seeking the minimum energy state and ③ extracting the solution from the final device state. In this subsection, we demonstrate how to map 3-SAT problems to an Ising Machine as a preliminary step.
@@
where each $\ell_{j,1}$ is a literal (either a variable $x_{p}$ or its negation $\overline{x}_{p}$) and each disjunction of literals $(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is a clause. The 3-SAT problem consists of finding some $x$ such that $f(x)=1$, i.e., an assignment satisfying all of the clauses. We restrict ourselves to 3-SAT without loss of generality, as any $K$-SAT problem can be reduced to an equivalent 3-SAT problem by adding $O(K\cdot m)$ variables and clauses [sipser_introduction_2012].
@@ II.3 Kuramoto Oscillator Model
where $\kappa$ and $J_{ij}$ describe global and pairwise interaction strengths respectively, and $\omega_{i}$ is the natural frequency of the $i^{th}$ oscillator. We set $\omega_{i}=0$ in Eq. 8 assuming that all oscillators have identical $\omega_{i}$ and viewing the system in a rotating reference frame. Using spin glass interaction coefficients for $J$ has led to oscillator-based accelerators for combinatorial optimization: oscillator Ising machines (OIM) [32, 4, 11, 24, 2, 7, graber2024firefly, sreedhara2024novel, garg2024phase, ochs2021ising, sreedhara2023digital, 31, 33].
@@
The coupled oscillator system tends to minimize energy $E(\theta)$ over time [lyapunov1992general], laying the foundation for OIMs that tackle combinatorial optimization problems. The original OIM architecture [32] utilized electronic LC-tank oscillators, and more recent proposals have broadened the choice of substrate to ring oscillators [20, 25], phase transition nano-oscillators [23, 22, 10], and magnetic tunnel junctions [14, 30].
@@ III.1 Hopf Oscillator Model
While the Kuramoto model is sufficient to study phase-only phenomena, real-world oscillators will have variability in both phase and amplitude. To enable a more faithful model, we can generalize from the phase-reduced representation to complex $z\in\mathbb{C}^{N}$. An oscillator model for combinatorial optimization and sampling is the Adronov-Hopf model (or “Hopf model” for brevity) [6, nikonov2015coupled]. The evolution of each Hopf oscillator is given by a system of coupled ODEs
@@ IV.1 Model Evaluation
Overall Performance: We evaluate model performance on 1,700 uniform random 3-SAT benchmarks from SATLIB [15]. Each set, ufN-M, denotes instances with N variables and M clauses; we use uf20-91, uf50-218, uf75-325, uf100-430, and uf150-645, all at a clause-to-variable ratio of approximately 4.3. This ratio lies at the empirical phase transition in satisfiability, where random 3-SAT instances shift from mostly satisfiable to mostly unsatisfiable and where difficulty peaks [monasson1999determining]. We define a solvable instance as one where at least one of 100 independent runs (with different initial conditions) reaches zero energy within a sufficient simulation time ($T=136$), allowing both models to achieve their best performance on the benchmark.
@@ ACKNOWLEDGMENTS
The data that support the findings of this article are openly available [ACAL_prapplied-oscillator-data].